OpenStudy (anonymous):

Enter the equation of the line meeting the given conditions. Please put the equation in standard form. Containing C(3, 3) and D(-3, 2)

OpenStudy (anonymous):

slope = 2-3/-3-(3) =-1/-6 =1/6 =M y-y1 = M(x-x1) y-3=1/6(x-3)